# Rethinking cluster-conditioned diffusion models

### v1

@@ 3.2 Cluster-conditional EDM (C-EDM) diffusion model
3.3 Estimating the upper cluster bound using TEMI clustering
We aim to find an upper cluster bound $C_{max}$ using TEMI ($\gamma=0.6$), such that $C_{V}<C_{max}$. During this computation, we neither make use of the generative model nor any additional information, such as $C_{GT}$, GT labels, or label names, in contrast to previous works [39, 7]. To this end, we denote the number of utilized clusters as $C^{u}$ after TEMI clustering. A cluster is utilized if at least one training sample is assigned to it. The TEMI cluster utilization ratio is defined as $r_{C}:=\frac{C^{u}}{C}\leq 1$.
Unlike $k$-means that always has $r_{C}=1$, we observe that as the number of clusters $C$ increases, the TEMI cluster utilization ratio $r_{C}$ typically decreases. This particularity is specific to NN pair-based clustering methods [87, 2], and there is no guarantee that the full cluster spectrum will be utilized. Intuitively, TEMI clustering with $\gamma=0.6$ is enforcing $r_{C}\rightarrow 1$, and the observation of $r_{C}<1$ is an indication that the maximum number of TEMI clusters is reached for $\gamma=0.6$. To obtain a simple estimate for $C_{max}$ based on the above intuition, we empirically define a cutoff threshold $r_{C}\leq 0.96$, based on a cross-dataset evaluation (Fig. 3).
Using binary search [50], starting from a small cluster value $C_{\text{start}}$, one can find ${C}_{max}$ in $O(\log_{2}(\frac{C_{max}}{C_{\text{start}}}))$ time, based solely on clustering iterations. Since there are many $C$ that can satisfy this criterion, we stop at the first one. Finally, if the number of GT labels is known, one can set $C_{\text{start}}=C_{GT}$. After detecting $C_{max}$, one can perform a hyperparameter search [93, 14] to find $C_{V}\in[2,C_{max})$.
@@ 4.1 Datasets, models, and metrics
Following prior works [39, 2, 96], we use DINO ViT-B (with a patch size of 16) pre-trained on ImageNet [28] for image clustering. We report the Fréchet inception distance (FID) [34] to quantify the image generation performance as it simultaneously captures visual fidelity and diversity. To facilitate future comparisons, we also compute the Fréchet DINOv2 distance (FDD) [83] that replaces the features of InceptionNet-v3 [85, 86] with DINOv2 ViT-L [63]. FID and FDD are averaged over three independently generated sample sets of 50K images each. To measure the cluster alignment w.r.t. GT labels, we use the adjusted normalized mutual information (ANMI) as in [40].
CIFAR10 CIFAR100 FFHQ-64
Generative methods Unlabelled Labelled Unlabelled Labelled Unlabelled
DDPM [35, 39] $3.17$ $10.7$ $9.7$
NCSN++-G [22] $2.25$ $3.86$
PFGM++ [90] - $1.74$ $2.43$
EDM [44] $1.97$ $1.79$ $2.39$
EDM (our reproduction) $2.07$ $1.81$ $3.41$ $2.21$ $2.53$
Self-conditional methods
IDDPM+$k$-means ($k$=10) [7] 2.23 - - - -
DDPM+$k$${}^{\dagger}$ ($k$=400) [39] - - $9.6$
C-EDM+$k$${}^{\dagger}$ (Ours) 1.69 - 2.21 - 1.99
${}^{\dagger}$ (Ours) 1.67 - 2.17 - $2.09$
[…]
@@
We follow the default hyperparameter setup from Karras et al. [44] for all our generative experiments and the setting of Adaloglou et al. [2] for TEMI clustering. We denote the number of million samples seen during training as $M_{img}$. We compare C-EDM with other baselines and state-of-the-art methods after $M_{img}=200$, while $M_{img}=100$ when comparing different cluster sizes. We train the TEMI clustering heads for 200 epochs per dataset. All the experiments were conducted on 4 NVIDIA A100 GPUs with 40GB VRAM each. On this hardware, TEMI clustering was more than 50×50$\times$50 × faster than training EDM on FFHQ-64, which requires more than $2$ days for 200 $M_{img}$ with a batch size of $512$. Additional implementation details and hyperparameters can be found in the supplementary.
We verify our approach across CIFAR10, CIFAR100 [51] and FFHQ. CIFAR10 and CIFAR100 have 50K samples and $32^{2}$ resolution images, while we use the 64x64 resized version of FFHQ (FFHQ-64) consisting of 70K samples. FFHQ illustrates the applicability of the upper bound introduced in Sec. 3.3 on unlabelled datasets. Intriguingly, we could visually identify TEMI clusters on FFHQ with easily distinguishable visual cues, such as groups of images with beanies, smiling faces, glasses, sunglasses, hats, and kids (see supplementary).
4.2 State-of-the-art and baseline comparison for image synthesis
We compare C-EDM with recent state-of-the-art generative models, including our reproduction of EDM using the official code in Tab. 1. Methods that focus on sampling-related modifications (e.g. classifier-free guidance) of diffusion models are not included, as they can be incorporated into any trained diffusion model [48, 74, 36]. With respect to FID, we found $C_{V}=100,200,400$ to be close to optimal cluster granularities for CIFAR10, CIFAR100, and FFHQ-64, respectively.
@@ 4.2 State-of-the-art and baseline comparison for image synthesis
Figure 2: FID score (y-axis) versus seen samples during training in millions ($M_{img}$, x-axis). TEMI and k-means clusters are computed using the representations of DINO ViT-B [18] model. We used $C_{V}=100,200,400$ for CIFAR10, CIFAR100 and FFHQ-64 respectively. The sample efficiency compared to the unconditional generation is indicated by the arrow. Best viewed in color.
@@
Compared to unconditional methods, C-EDM achieves an average relative FID improvement of 24.4% and 24.9% using TEMI and k-means clusters, respectively. We also report small improvements compared to GT label conditioning on CIFAR10 and CIFAR100 (i.e. 4.77% mean relative improvement in FID using TEMI), which showcases the effectiveness of C-EDM. More importantly, previous cluster-conditional approaches did not achieve near-state-of-the-art FIDs because they: a) adopted non-competitive diffusion baselines such as DDPM [35], b) did not consider pre-trained feature extractors for clustering [7], c) did not use the optimal cluster granularity [39]. For instance, Hu et al. [39] used $400$ clusters on CIFAR100 using DINO ViT-B for clustering, while Fig. 3 shows that $200$ clusters lead to a superior FID using C-EDM (11.2% relative improvement).
@@
Sample efficiency. In addition to the reported gains in FID, we study the sample efficiency of C-EDM across datasets in Fig. 2. On CIFAR10 and CIFAR100, the sample efficiency for $C=C_{V}$ compared to the unconditional model ($C=1$) peaks at 5×5$\times$5 ×, where C-EDM with $M_{img}=40$ outperforms the unconditional model at $M_{img}=200$. On FFHQ-64, which is not a classification dataset, we report a sample efficiency of $\times3$. TEMI and $k$-means demonstrate identical sample efficiency compared to the unconditional model. Compared to GT conditional EDM, we find that the sample efficiency of C-EDM against GT conditional EDM heavily depends on the granularity of the GT labels. We suspect the smaller the gap between $C_{GT}$ and $C_{V}$, the smaller the sample efficiency.
@@
Figure 3: We report FID (left y-axis) and TEMI cluster utilization ratio $r_{C}$ (right y-axis) across different clusters using C-EDM, evaluated at $M_{img}=100$. The green area indicates the discovered upper cluster bound for $r_{C}\leq 0.96$ using only clustering trials.
@@ 4.3 Cluster utilization ratio and discovered upper bounds
We depict how the TEMI cluster utilization $r_{C}$ changes in tandem with FID for different cluster sizes in Fig. 3. Across three datasets, we identify that for $r_{C}\leq 0.96$, FID deteriorates significantly, which we use as an upper bound in this work. Crucially, this approach also applies to FFHQ, where no GT labels exist. Moreover, we can discern certain patterns across datasets: (i) $C_{V}$ always has a high utilization ratio $r_{C_{V}}>0.96$, (ii) the majority of experiments with $C<C_{max}$ outperform the unconditional model at $M_{img}=100$, and (iii) $C_{V}>C_{GT}$ for CIFAR10 and CIFAR100, which is in line with [39]. Generated FFHQ-64 samples corresponding to various cluster sizes are depicted in Fig. 4.
Computational complexity comparison. We emphasize that the introduced upper bound can be computed by measuring $r_{C}$ without GT labels or training the generative model, which on large scales (ImageNet) can require up to $4$MWh per experiment [46]. On the considered hardware, we find that TEMI clustering was 50x faster on FFHQ-64 and estimate a speedup of $\approx 400$ on ImageNet at $64^{2}$ resolution with the default setup of Karras et al. [44].
Figure 4: Generated FFHQ-64 samples using C-EDM and TEMI clusters with different granularity levels $C$ as well as unconditional EDM (Uncond.). All samples in a row use the same initial noise. The cluster assignment is randomly sampled for each $C$.
@@ 4.4 Investigating the connection between clustering and cluster-conditional image synthesis
Table 2: FID for different grouping methods as a condition to EDM ($M_{img}=100$). We use the 20 GT superclasses and 100 GT labels of CIFAR100. CLIP pseudo-labels are computed using zero-shot classification based on the GT label names.
CIFAR100
$C$ 20 100 200
Human annotation 3.10 2.41 -
@@
[Uncaptioned image] Figure 5: FID of C-EDM (y-axis) across different cluster sizes $C$ (x-axis) for multiple feature extractors on CIFAR10 evaluated at $M_{img}=100$. The ANMI is shown in parentheses for $C_{V}$=100.
Table 2: FID for different grouping methods as a condition to EDM ($M_{img}=100$). We use the 20 GT superclasses and 100 GT labels of CIFAR100. CLIP pseudo-labels are computed using zero-shot classification based on the GT label names.
C-EDM matches EDM when $C=C_{GT}$. In Fig. 3, we observe that the GT label conditioning closely follows the FID of TEMI clusters for $C=C_{GT}$. In Tab. 2, we leverage the two annotation levels of CIFAR100 [51], specifically the 20 GT superclasses and the 100 GT labels to benchmark how different grouping methods perform in image synthesis. Apart from image clustering, we create text-based pseudo-labels with CLIP (OpenCLIP ViT-G/14 [41]) using prompts such as “a photo of a {label}” similar to [60]. Then, CLIP pseudo-labels are derived from zero-shot classification [69] using the $\operatorname*{arg\,max}$ of the text-image similarity after softmax [1]. For reference, we provide the FID with the number of visual groups $C_{V}=200$ for $k$-means and TEMI clusters. Interestingly, all methods attain a similar FID for $C_{GT}=100$, while TEMI achieves the best FID only for $C=20$. This suggests that generative performance is mostly invariant to the grouping method, including human annotation, given a pre-trained backbone that captures general visual concepts. Next, we investigate the impact of the choice of pre-trained backbone.
The number of visual groups is independent of the backbone. The broader adoption of the discovered $C_{V}$ for each dataset requires our analysis to be independent of the backbone. In Fig. 5, we show that all pre-trained backbones achieve similar FIDs across different cluster sizes even though the ANMI varies up to 5.4% on CIFAR10. The backbones differ in terms of the number of parameters (86M up to 1.8 billion), pre-training dataset size (1.2M up to 2B), and objective (MoCO[24], DINO[18], iBOT[95], CLIP[41, 69]). It is worth noting that the representations of MoCOv3 have the worst ANMI, while its clusters lead to a competitive FID, which is consistent with concurrent work [55]. The fact that differences in clustering performance between backbones do not translate to improvements in FID is in contrast with Hu et al. [39].
@@
Table 3: Effect of self-supervised fine-tuning (adapted weights) [72] on the pretrained DINO ViT-B (ImageNet weights). TEMI is used for clustering. ACC stands for the 20-NN classification accuracy using the backbone features. We measure FID and FDD for both models at $M_{img}=100$.
@@ 5 Discussion
The clusters’ granularity level determines conditional generative performance. In general, the quality of off-the-shelf feature representations is typically determined by how separable the images are in the feature space w.r.t. GT labels [64]. However, our experimental analysis shows that in image synthesis, the quality of image-level conditioning primarily depends on the granularity of the cluster assignments. Overall, we find no significant connection between the task of feature-based clustering and cluster-conditional image generation. The pre-trained backbone and the categorization method do not severely affect generative performance for the considered datasets. Unlike human annotation, which is inefficient, costly, and prone to errors, clustering can create groups of arbitrary size suitable to the application.
@@
Figure 6: Number of clusters (x-axis) versus the average FID (y-axis) for CIFAR10 (left), CIFAR100 (center), and FFHQ-64 (right) using C-EDM after $M_{img}=100$. The non-filled markers indicate unconditional EDM ($C=1$). Best viewed in color.
@@
The optimal number of visual groups depends on the generative metric. So far, we have used the commonly adopted FID score to select $C_{V}$. Here, we investigate the impact of the choice of generative metric by comparing FDD and FID. We highlight that the only difference is that FID and FDD use InceptionV3 and DINOv2 features, respectively. The cross-dataset evaluation in Fig. 6 shows that FID and FDD do not always agree with respect to the number of visual groups. The largest disagreement is observed on CIFAR10. Still, the discovered upper bound $C_{max}$ always includes the number of visual groups with respect to FID and FDD. Finally, we evaluate FID and FDD across training iterations and found that while FID fluctuates after $M_{img}=120$, FDD decreases monotonically (see Supplementary Material).
Apart from FID and FDD, we measured additional feature-based metrics such as precision, recall, diversity, coverage, and mean similarity score (MSS [74]) using DINOv2 [83] and could not identify any pattern across clusters. The Inception Score (IS) showed negligible differences across clusters and cannot be used for arbitrary unsupervised datasets since the class overlap with ImageNet cannot be guaranteed. Similar to [26, 65, 42, 83], we argue that new generative metrics must be developed, especially regarding diversity and coverage.

### v2

@@ 3.2 Cluster-conditional EDM (C-EDM)
CIFAR10 CIFAR100 FFHQ-64 FFHQ-128
Generative methods Unlabelled GT Unlabelled GT Unlabelled Unlabelled
DDPM [29, 34] $3.17$ $10.7$ $9.7$
PFGM++ [79] - $1.74$ $2.43$
EDM [39] $1.97$ $1.79$ $2.39$
EDM (our reproduction) $2.07$ $1.81$ $3.41$ $2.21$ $2.53$ 5.93
Self-conditional methods
IDDPM+$k$-means ($k$=10) [25] 2.23 - - - - -
DDPM+$k$-means ($k$=400) [34] - - $9.6$
C-EDM+$k$-means (Ours) 1.69 - 2.21 - 1.99 -
C-EDM+TEMI (Ours) 1.67 - 2.17 - $2.09$ 4.40
EDM+oracle features (Ours) 2.21 - 2.25 - 1.77 -
Table 1: State-of-the-art generative model comparison: FID ($\downarrow$) for various GT labeled and unlabelled benchmarks. We use $C_{V}=$100, 200, 400 clusters for CIFAR10, CIFAR100, and FFHQ-64, respectively. Self-conditional methods with use the pre-trained DINO ViT-B feature extractor. Ground truth and oracle feature conditioning results are marked in gray color as they are not a fair comparison.
3.3 Estimating the upper cluster bound
We aim to find an upper cluster bound $C_{max}$ using TEMI ($\gamma=0.6$), such that $C_{V}<C_{max}$. During this computation, we neither use the generative model nor any additional information. We highlight that previous works [50, 25, 34] have overlooked this design choice and iterate within a small range around $C_{GT}$. We denote the number of utilized clusters (one training sample is assigned to it) as $C^{u}$ after TEMI clustering. The TEMI cluster utilization ratio is defined as $r_{C}:=\frac{C^{u}}{C}\leq 1$. Importantly, there is no guarantee that the full cluster spectrum will be utilized. Unlike $k$-means that always has $r_{C}=1$, we observe that as the number of clusters $C$ increases, the TEMI cluster utilization ratio $r_{C}$ typically decreases, which provides dataset-specific information. Intuitively, TEMI clustering with $\gamma=0.6$ is enforcing $r_{C}\rightarrow 1$, and the observation of $r_{C}<1$ is an indication that the maximum number of TEMI clusters is reached for $\gamma=0.6$.
The search for $C_{max}$ involves doubling the number of clusters $C$ at each iteration until the $r_{C}$ falls below a threshold, $r_{C}\leq\alpha$, followed by a more fine-grained grid search. $C_{max}$ is then defined as the highest $C$ for which $r_{C}>\alpha$. This results in a worst-case time complexity of $O(\log_{2}(\frac{C_{max}}{C_{\text{start}}}))$. This is conceptually similar to the elbow method [75, 42]. In this sense, the proposed heuristic has the same limitations as the elbow method. Yet, it is currently the only method that provides a practitioner with a starting point for an unlabelled dataset. After detecting $C_{max}$, we perform a bounded grid search to find $C_{V}\in[2,C_{max})$ using the generative model. To obtain a simple cross-dataset estimate, we empirically find a cutoff threshold $r_{C}\leq 0.96$ to work well across three datasets.
@@ 4.1 Datasets, models, and metrics
Following prior works [34, 2, 83], we use DINO ViT-B pre-trained on ImageNet [22] for image clustering. We report the Fréchet inception distance (FID) [28] to quantify the image generation performance as it simultaneously captures visual fidelity and diversity. To facilitate future comparisons, we also compute the Fréchet DINOv2 distance (FDD) [71] that replaces the features of InceptionNet-v3 [73, 74] with DINOv2 ViT-L [54]. FID and FDD are averaged over three independently generated sample sets of 50K images each. To measure the cluster alignment w.r.t. GT labels, we use the adjusted normalized mutual information (ANMI) as in [35].
@@
We follow the default hyperparameter setup for EDM [39] and TEMI [2]. We denote the number of samples (in millions) seen during training as $M_{img}$. We use $M_{img}=200$ when comparing C-EDM with other baselines and state-of-the-art methods (Tab. 1), and $M_{img}=100$ when comparing across different number of clusters. We train the TEMI clustering heads for 200 epochs per dataset. All the experiments were conducted on 4 NVIDIA A100 GPUs with 40GB VRAM each. On this hardware, TEMI clustering was more than 50×50$\times$50 × faster than training EDM on FFHQ-64, which requires more than $2$ days for 200 $M_{img}$ with a batch size of $512$. Additional implementation details and hyperparameters can be found in the supplementary material. We verify our approach across CIFAR10, CIFAR100 [45] and FFHQ. CIFAR10 and CIFAR100 have 50K samples and $32^{2}$ resolution images, while we use the 64x64 and 128x128 versions of FFHQ (FFHQ-64, FFHQ-128) consisting of 70K samples. Finally, diffusion sampling methods are not included, as they can be incorporated into any diffusion model [43, 64, 30].
@@
Figure 2: FID (y-axis) versus seen samples during training in millions (x-axis). TEMI and k-means clusters are computed using the representations of DINO ViT-B [15]. We used $C_{V}=100,200,400$ for CIFAR10, CIFAR100 and FFHQ-64 respectively. The training sample efficiency compared to the unconditional baseline is indicated by the arrow. Best viewed in color.
4.2 State-of-the-art comparison for image synthesis
@@
Comparison with state-of-the-art unconditional generative models. With respect to FID, we found $C_{V}=100,200,400$ to be close to optimal cluster granularities for CIFAR10, CIFAR100, and FFHQ-64, respectively. We use these values to report FIDs compared to GT conditional and unconditional generative models in Tab. 1. Compared to unconditional methods, C-EDM achieves an average relative FID improvement of 24.4% and 24.9% using TEMI and k-means clusters, respectively. More importantly, previous cluster-conditional approaches did not achieve near-state-of-the-art FIDs because they: a) adopted non-competitive diffusion baselines such as DDPM [29], b) did not consider pre-trained feature extractors for clustering [25], c) did not use the optimal cluster granularity [34]. For instance, Hu et al. [34] used $400$ clusters on CIFAR100 using DINO ViT-B for clustering, while Fig. 3 shows that $200$ clusters lead to a superior FID using C-EDM (11.2% relative improvement).
Comparison with state-of-the-art GT conditional generative models. Intriguingly, using C-EDM with $C_{V}$ clusters, we report small improvements compared to GT label conditioning on CIFAR10 and CIFAR100 in Tab. 1. For instance, 4.77% mean relative improvement in FID using TEMI. Even though cluster-conditioning is primarily designed for unconditional generation, such as FFHQ, we demonstrate that it can match or outperform GT labels, which showcases the effectiveness of C-EDM.
@@
Sample efficiency. In addition to the reported gains in FID, we study the sample efficiency of C-EDM during training across datasets in Fig. 2. On CIFAR10 and CIFAR100, the training sample efficiency for $C=C_{V}$ compared to the unconditional model ($C=1$) peaks at 5×5$\times$5 ×, where C-EDM with $M_{img}=40$ outperforms the unconditional model at $M_{img}=200$. On FFHQ-64, which is not a classification dataset, we report a sample efficiency of $\times3$. TEMI and $k$-means demonstrate identical sample efficiency compared to the unconditional model. Intuitivly, a more informative conditioning signal enables learning the data distribution faster. Upon visual inspection, we could identify FFHQ clusters for $C=C_{V}$ with easily distinguishable visual patterns, such as groups of images with beanies, smiling faces, glasses, sunglasses, hats, and kids (see supplementary). Finally, we find that the sample efficiency of C-EDM against GT conditional EDM heavily depends on the quality and granularity of the GT labels.
@@
Figure 3: FID (left y-axis) and TEMI cluster utilization ratio $r_{C}$ (right y-axis) across different numbers of clusters $C$ (x-axis) using C-EDM, evaluated at $M_{img}=100$. The green area indicates the discovered cluster range $[2,C_{max})$ for $r_{C}\leq\alpha=0.96$.
@@ 4.3 Cluster utilization ratio and discovered upper bounds
We depict how the TEMI cluster utilization $r_{C}$ changes in tandem with FID for different numbers of clusters in Fig. 3. Crucially, this approach also applies to FFHQ, where no GT labels exist. Moreover, we can discern certain patterns across datasets: (i) $C_{V}$ always has a high utilization ratio, (ii) the majority of experiments with $C<C_{max}$ outperform the unconditional model at $M_{img}=100$, and (iii) $C_{V}>C_{GT}$ for CIFAR10 and CIFAR100, which is in line with [34]. Even though the choice of $\alpha=0.96$ is not guaranteed to be generally applicable and is based on empirical evidence, a more or less strict choice can be used based on the practitioner’s computational budget. The introduced upper bound can be computed with negligible computational overhead and without access to GT labels or training the generative model, which on large scales (ImageNet) can require up to $4$MWh per experiment [40].
@@ 4.4 Investigating the connection between clustering and cluster-conditional image synthesis
$C$
CIFAR100 20 100 200
Human annotation (GT) 3.10 2.41 -
@@
Table 2: FID for different grouping methods as a condition to EDM ($M_{img}=100$). We use the 20 GT superclasses and 100 GT labels of CIFAR100. CLIP pseudo-labels are computed using zero-shot classification pseudo-labels based on the GT label names.
C-EDM matches EDM when $C=C_{GT}$. In Fig. 3, we observe that the GT label conditioning closely follows the FID of TEMI clusters for $C=C_{GT}$. In Tab. 2, we leverage the two annotation levels of CIFAR100 [45], specifically the 20 GT superclasses and the 100 GT labels to benchmark how different grouping methods perform in image synthesis. Apart from image clustering, we create text-based pseudo-labels with CLIP (OpenCLIP ViT-G/14 [36]) similar to [51]. Then, CLIP pseudo-labels are derived from zero-shot classification [59] using the $\operatorname*{arg\,max}$ of the text-image similarity after softmax [1]. For reference, we provide the FID with the number of visual groups $C_{V}=200$ for $k$-means and TEMI clusters. Interestingly, all methods attain a similar FID for $C_{GT}=100$, while TEMI achieves the best FID only for $C=20$. This suggests that generative performance is mostly invariant to the grouping method, including human annotation, given a feature extractor that captures general visual concepts. Next, we investigate the impact of the pre-trained feature extractor.
Figure 4: FID (y-axis) across different numbers of clusters $C$ (x-axis) using C-EDM with TEMI with different feature extractors. The ANMI is shown in parentheses for $C_{V}$=100.
@@
Table 3: Effect of self-supervised fine-tuning (adapted weights) on the pre-trained DINO ViT-B. ACC stands for the 20-NN classification accuracy using the DINO features. We measure FID and FDD for both models at $M_{img}=100$.
The number of visual groups is insensitive to the chosen feature extractor. The broader adoption of the discovered $C_{V}$ for each dataset requires our analysis to be insensitive to the chosen feature extractor. In Fig. 4, we show that the pre-trained models considered in this work achieve similar FIDs across different numbers of clusters even though the ANMI varies up to 5.4% on CIFAR10. Feature extractors differ in terms of the number of parameters (86M up to 1.8 billion), pre-training dataset size (1.2M up to 2B), and objective (MoCO[19], DINO[15], iBOT[82], CLIP [36, 59]). Our finding is consistent with concurrent work on the scale of ImageNet [49] that shows that linear probing accuracy is not indicative of generative performance for feature-based conditioning. The fact that differences in discriminative performance do not translate to improvements in FID is in contrast with Hu et al. [34].
@@
Figure 5: Top 1-NN cosine similarity AUROC (left y-axis) and Frechet distance between the C-EDM and unconditional samples (uFID) for different cluster sizes $C$ (x-axis). For the computation of AUROC, we use the official test splits.
@@ 5 Discussion
Highly fine-grained clusters lead to out-of-distribution samples. To quantify the degree of out-of-distribution for the generated C-EDM samples, we measure the AUROC using the top-1 NN cosine similarity (1-NN) using DINO ViT-B [72] and compare with the test split of CIFAR10 and CIFAR100 (Fig. 5). We note that 1-NN is independent of the data distribution of the generated samples and thus does not account for diversity. The highest $C$ considered for CIFAR10 and CIFAR100 produces the highest AUROC, suggesting that the features of the generated samples are pushed away from the training data. We hypothesize that this behavior originates from highly specialized clusters that cannot always be generated from the initial noise, resulting in visual artifacts. In parallel, we measure how similar the C-EDM samples are compared to the unconditional ones using FID. We call this metric unconditional FID (uFID). We observe that uFID increases as $C$ increases, suggesting that the generated samples for larger $C$ are farther away in feature space than the unconditional samples.
@@
The clusters’ granularity level determines conditional generative performance. The quality of feature representations is typically determined by linear separability w.r.t. GT labels [55]. However, our experimental analysis shows that in image synthesis, the quality of image-level conditioning primarily depends on the granularity of the cluster assignments (Tab. 2 and Fig. 3). Overall, we find no significant connection between the task of feature-based clustering and cluster-conditional image generation (Fig. 4 and Tab. 3). The pre-trained model and the categorization method do not severely affect generative performance for the considered datasets.
@@
Figure 6: FID (left y-axis) and FDD (right y-axis) versus number of clusters (x-axis) using C-EDM at $M_{img}=100$. The non-filled markers indicate the unconditional EDM.
@@
Generative metric and optimal number of clusters. Here, we investigate the impact of the choice of generative metric by comparing FDD and FID. We highlight that the only difference is that FID and FDD use InceptionV3 and DINOv2 features, respectively. The cross-dataset evaluation in Fig. 6 shows that FID and FDD do not always agree with respect to the number of visual groups. The largest disagreement is observed on CIFAR10. Still, the discovered upper bound $C_{max}$ always includes the number of visual groups. Moreover, we evaluate both metrics across training iterations and found that while FID fluctuates after $M_{img}=120$, FDD decreases monotonically (supplementary material). Finally, we measured additional feature-based metrics such as precision, recall, diversity, coverage, MSS, and inception score without success. Similar to [21, 47, 56, 37, 71], we argue that new generative metrics must be developed.
Figure 7: Comparing C-EDM samples generated with learned cluster assignments (top row, C=100) to EDM samples generated with GT labels, on CIFAR100. The clusters are mapped to GT classes using the Hungarian algorithm.
Visual comparison of C-EDM to EDM. In Fig. 7, we map clusters to their respective CIFAR100 classes and produce samples with C-EDM and GT conditional EDM using the same noise. Samples exhibit high visual similarity, corroborating with Tab. 2 and Fig. 2. More visualizations are provided in the supplementary material.
Measuring the confidence of the generated samples using TEMI. In Fig. 8, we show the generated examples with the lowest and highest maximum softmax probability [27] of the TEMI head as a measure of confidence. For comparison, we show unconditional samples generated using the same initial noise in the denoising process. Visual inspection shows that low-confidence C-EDM samples do not have coherent semantics compared to the unconditional ones, leading to inferior image quality. We hypothesize that the sampled condition for the low-confidence samples is in conflict with the existing patterns in the initial noise. Increasing the number of conditions likely leads to worse image-condition alignment [65, 30, 32]. We argue that internal guidance methods could be employed to increase the image quality in such cases, which is left for future work.
(a) High-confidence C-EDM samples
(b) Low-confidence C-EDM samples
Figure 8: Generated high (a) and low (b) confidence CIFAR10 samples. The top row depicts the unconditional (Uncond.) samples, while the bottom row shows the generated samples using C-EDM with TEMI ($C=100$). Images on the same column are produced with the same initial noise.
By contrast, highly confident C-EDM samples show more clearly defined semantics than unconditional ones. When the low frequencies, such as the object’s shape, remain intact, cluster conditioning aids in refining local pixel patterns. Confident C-EDM samples consist of simple pixel patterns that are easy to generate, such as a white background. Confidence can also be leveraged in future works in rejection sampling schemes [17]. More samples are provided in the supplementary material.
6 Conclusion
In this paper, a systematic empirical study was conducted focusing on conditioning diffusion models with cluster assignments. It was demonstrated that cluster conditioning achieves state-of-the-art FID on three generative benchmarks while attaining strong sample efficiency. To reduce the search space for estimating the visual groups of the dataset, a novel method that computes an upper cluster bound based solely on clustering was proposed. Finally, our experimental study indicates that generative performance using cluster assignments depends primarily on the granularity of the assignments.
Appendix A Visualizations
In Fig. 10, we map the TEMI clusters to classes using the Hungarian mapping [46] on CIFAR100. For the mapping to be one-to-one, we set $C=100$. We then generate samples using the same initial noise with both C-EDM and GT-conditioned EDM and visualize the first 20 clusters on CIFAR100. Given the same initial noise and the cluster that is mapped to its respective GT class, we observe a lot of visual similarities in the images, even though the two models (C-EDM and EDM) have different weights and have been trained with different types of conditioning.
In Figs. 11 and 12, we visualize C-EDM samples generated from the same initial noise on FFHQ-64 for diffusion models trained with varying cluster granularities. Each noise gets a condition sampled from $p(c)$. Similar to our quantitative analysis, the generated images from small cluster sizes are closer to the unconditional prediction. Finally, in Fig. 9, we visualize cluster-conditioned and unconditional FFHQ samples at $M_{img}=100M$.
Figure 9: Visual comparison between C-EDM (C=400) and unconditional EDM (Uncond.) at $M_{img}=100M$ on FFHQ at 128x128.
In Fig. 13, we visualize real training FFHQ images that are grouped in the same TEMI cluster using the DINO features. We visually identify groups with shared characteristics such as sunglasses, hats, beanies, pictures of infants, and pictures of speakers talking to a microphone. Finally, in Fig. 14 we provide a more detailed visual comparison of low and high confidence samples using C-EDM on CIFAR100.
Figure 10: Visualizing generated images from CIFAR100 using C-EDM (even rows) and ground truth conditional EDM (odd rows) using the same initial noise and deterministic noise sampling. We map the C=100 CIFAR100 cluster to the respective ground truth class as computed via the Hungarian one-to-one mapping.
Figure 11: Visualizing generated images from FFHQ-64 using CEDM for different number of clusters $C$ with the same random noise. We use deterministic noise sampling. Each noise gets a condition sampled from $p(c)$ for each individual clusters.
Figure 12: Generated FFHQ-64 samples using C-EDM and TEMI clusters with different granularity levels $C$ as well as unconditional EDM (Uncond., first column). All samples in a row use the same initial noise. The cluster assignment is randomly sampled from $q(c)$ for each $C$.